# Quantitative periodic homogenization of parabolic equations with large drift and potential

### v2

@@ Acknowledgements
The author would like to thank his PhD advisor Harsha Hutridurga for introducing him to the exciting theory of homogenization and for his useful remarks during the preparation of this manuscript. The author has also greatly benefited from the discussions with Chandan Biswas and Debanjana Mitra on a preliminary version of this manuscript. The author acknowledges the support of NBHM-DAE (Government of India).
2 The case of merely bounded coefficients
@@
$\left\{\begin{aligned} \zeta\left(\frac{x}{\varepsilon}\right)\frac{\partial f_{\varepsilon}}{\partial t}-\operatorname{div}\left(\Theta\left(\frac{x}{\varepsilon}\right)\nabla f_{\varepsilon}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ f_{\varepsilon}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ f_{\varepsilon}(0,x)&=f_{\text{in}}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (2.1)
@@
$\zeta(y)\geq c>0\,\mbox{ for a.e. }y\in Y\quad\mbox{ and }\quad\int_{Y}\zeta(y)\,{\rm d}y=1,$ (2.2)
@@
$\left\{\begin{aligned} \operatorname{div}_{y}\left(\Theta(y)\nabla_{y}\left(\omega_{j}(y)+y_{j}\right)\right)&=0\qquad\mbox{ in }Y,\\ \int_{Y}\zeta(y)\omega_{j}(y)\,{\rm d}y&=0,\end{aligned}\right$. (2.3)
@@
$\left\{\begin{aligned} \frac{\partial f_{0}}{\partial t}-\operatorname{div}\left(\Theta_{h}\nabla f_{0}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ f_{0}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ f_{0}(0,x)&=f_{\text{in}}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (2.4)
@@
$\left\|f_{\varepsilon}-f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}.$ (2.5)
@@
Next, we consider a parabolic smoothing operator borrowed from [12, Section 3, p.2102]:
$S_{\varepsilon}(g)(t,x):=\frac{1}{\varepsilon^{d+2}}\int_{\mathbb{R}^{d+1}}g(t-s,x-y)\,\theta\left(\frac{s}{\varepsilon^{2}},\frac{y}{\varepsilon}\right)\,{\rm d}y\,{\rm d}s,$ (2.6)
@@
$w_{\varepsilon}:=f_{\varepsilon}-f_{0}-\varepsilon\omega\left(\frac{x}{\varepsilon}\right)\cdot S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right),$ (2.7)
where $f_{\varepsilon}$ and $f_{0}$ are solutions to (2.1) and (2.4) respectively. Note that $w_{\varepsilon}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$ and that $w_{\varepsilon}(0,\cdot)\equiv 0$. The study of a function defined similar to (2.7) is carried out in the quantitative analysis of periodic elliptic problems [14, Chapter 3, page 33]. We note, however, that the analysis carried out by Geng and Shen in [12, equation (1.14), page 2095] considers a function which also involves a term of order $\mathcal{O}(\varepsilon^{2})$. We will be employing such a construction in the next section. In the remainder of this manuscript, we will sometimes be using the following shorthand notations:
@@
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$ (2.8)
@@
$\operatorname{div}_{y}\phi(y)=\Theta(y)+\Theta(y)\nabla_{y}\omega(y)-\Theta_{h}\qquad\mbox{ in }Y.$ (2.9)
@@
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$ (2.10)
@@
where we have used the fact that $f_{\varepsilon}$ and $f_{0}$ satisfy the evolution equations in (2.1) and (2.4), respectively. Owing to the anti-symmetric nature of the third-order tensor (i.e. $\phi_{kij}=-\phi_{ikj}$ for $1\leq i,j,k\leq d$), it follows that (see [14, Remark 3.1.3]) for a nice function $\psi$, we have
@@
$\left|\int_{\Omega}\tau\left(\frac{x}{\varepsilon}\right)\kappa(x)\,{\rm d}x\right|\leq C\varepsilon\left\|\nabla\kappa\right\|_{\mathrm{L}^{\frac{2d}{d+2}}(\Omega)}.$ (2.11)
@@
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi$ (2.12)
@@
where we have used the Lemma 4.1 from the Appendix. Note that
@@
where we have used Lemma 4.2 from the Appendix and have employed the Hölder inequality in the last step. Using a property of $S_{\varepsilon}$ (see Lemma 4.3 from the Appendix), we arrive at the following bound:
@@
where we have used the fact [12, Remark 3.6, page 2106] that for any $g\in\mathrm{H}^{1}(\Omega)$,
@@
and the following estimate [12, Remark 3.10, page 2108]:
@@
$\nabla\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}=S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)+S_{\varepsilon}\Big(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla^{2}f_{0}\Big).$ (2.13)
@@
where we have used Lemma 4.1 from the Appendix. Using the property (2.13), and arguing as we did in the case of $I_{4}$, we can arrive at the following bound for $I_{5}$:
@@
3 The case of regular coefficients
@@
$\left\|f_{\varepsilon}-f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{2}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}.$ (3.1)
@@
$P:=\Theta(y)+\Theta(y)\nabla_{y}\omega(y)-\Theta_{h}\qquad\mbox{ in }Y.$ (3.2)
@@
$H_{ij}:=\left\{\begin{aligned} P_{ij}&\qquad\mbox{ for }1\leq i,j\leq d,\\ -\zeta(y)\omega_{j}(y)&\qquad\mbox{ for }i=d+1\mbox{ and }1\leq j\leq d.\end{aligned}\right$. (3.3)
@@
Hence there exists a third order tensor $\mathfrak{B}$ (see [14, Proposition 3.1.1, page 35]) such that
$\frac{\partial\mathfrak{B}_{(d+1)ij}}{\partial\tau}+\sum_{k=1}^{d}\frac{\partial\mathfrak{B}_{kij}}{\partial y_{k}}=H_{ij}\qquad\mbox{ for }1\leq i\leq d+1,\quad 1\leq j\leq d,$ (3.4)
@@
$\mathfrak{B}_{kij}=-\mathfrak{B}_{ikj}\qquad\mbox{ for }1\leq i,k\leq d+1,\qquad 1\leq j\leq d.$ (3.5)
@@
$z_{\varepsilon}:=f_{\varepsilon}-f_{0}-\varepsilon\omega\left(\frac{x}{\varepsilon}\right)\cdot S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\nabla f_{0}\right)-\varepsilon^{2}\sum_{i,j=1}^{d}\frac{\mathfrak{B}_{(d+1)ij}\left(\frac{x}{\varepsilon}\right)}{\zeta\left(\frac{x}{\varepsilon}\right)}\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right),$ (3.6)
@@
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}z_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla z_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$ (3.7)
@@
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}z_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla z_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$ (3.8)
@@
$\ell_{1}\leq C\varepsilon\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;\mathrm{W}^{1,d}(\Omega))}\left\|\psi\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}.$ (3.9)
@@
$\ell_{2}\leq C\varepsilon^{\frac{1}{2}}\left\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{\text{in}}\right\|_{\mathrm{L}^{2}(\Omega)}\right\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$ (3.10)
@@
$\ell_{3}\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$ (3.11)
@@
$\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right)=S_{\varepsilon}\left(\frac{\partial}{\partial x_{i}}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\frac{\partial^{2}f_{0}}{\partial x_{i}\partial x_{j}}\right)$ (3.12)
@@
$\nabla^{2}S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\nabla f_{0}\right)=\nabla S_{\varepsilon}\left(\nabla\left(\eta^{\varepsilon}_{3}\eta^{\varepsilon}_{4}\right)\nabla f_{0}\right)+\nabla S_{\varepsilon}\left(\eta^{\varepsilon}_{3}\eta^{\varepsilon}_{4}\nabla^{2}f_{0}\right)$ (3.13)
@@
Using (3.12) and (3.13) and a few properties of the smoothing operator $S_{\varepsilon}$ (see Lemma 4.1 and 4.3 in the Appendix), we obtain the bound:
$\ell_{4}$ $\leq C\|\nabla f_{0}\|_{\mathrm{L}^{2}(\Omega_{T,3\varepsilon})}\|\nabla\psi\|_{\mathrm{L}^{2}(\Omega_{T,3\varepsilon})}+C\varepsilon\|\nabla^{2}f_{0}\|_{\mathrm{L}^{2}(\Omega_{T})}\|\nabla\psi\|_{\mathrm{L}^{2}(\Omega_{T})}$ (3.14)
@@
Note that to arrive at the estimate (3.14), we have used the fact that $\frac{\mathfrak{B}_{(d+1)ij}}{\zeta}\in\mathrm{H}^{1}(\Omega)$ – see Lemma 4.4 in the Appendix. We emphasise here that it is at this point we require the coefficient $\zeta$ to be more regular. The terms $\ell_{5}$ and $\ell_{6}$ are similar to $I_{5}$ in the proof of Lemma 2.4 (again, except for the choice of the cut-off functions) and hence can be handled in a similar fashion while taking into account the identity:

### v3

@@ Acknowledgements
The author would like to thank his PhD advisor Harsha Hutridurga for introducing him to the exciting theory of homogenization and for his useful remarks during the preparation of this manuscript. The author has greatly benefited from the discussions with Chandan Biswas and Debanjana Mitra on a preliminary version of this manuscript. The author would also like to thank Grégoire Allaire for pointing him to the reference [5]. Finally, the author acknowledges the support of NBHM-DAE (Government of India).
2. The case of merely bounded coefficients
@@
(2.1) $\left\{\begin{aligned} \zeta\left(\frac{x}{\varepsilon}\right)\frac{\partial f_{\varepsilon}}{\partial t}-\operatorname{div}\left(\Theta\left(\frac{x}{\varepsilon}\right)\nabla f_{\varepsilon}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ f_{\varepsilon}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ f_{\varepsilon}(0,x)&=f_{\text{in}}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$.
@@
(2.2) $\zeta(y)\geq c>0\,\mbox{ for a.e. }y\in Y\quad\mbox{ and }\quad\int_{Y}\zeta(y)\,{\rm d}y=1,$
@@
(2.3) $\left\{\begin{aligned} \operatorname{div}_{y}\left(\Theta(y)\nabla_{y}\left(\omega_{j}(y)+y_{j}\right)\right)&=0\qquad\mbox{ in }Y,\\ \int_{Y}\zeta(y)\omega_{j}(y)\,{\rm d}y&=0,\end{aligned}\right$.
@@
(2.4) $\left\{\begin{aligned} \frac{\partial f_{0}}{\partial t}-\operatorname{div}\left(\Theta_{h}\nabla f_{0}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ f_{0}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ f_{0}(0,x)&=f_{\text{in}}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$.
@@
(2.5) $\left\|f_{\varepsilon}-f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}.$
@@
Next, we consider a parabolic smoothing operator borrowed from [14, Section 3, p.2102]:
(2.6) $S_{\varepsilon}(g)(t,x):=\frac{1}{\varepsilon^{d+2}}\int_{\mathbb{R}^{d+1}}g(t-s,x-y)\,\theta\left(\frac{s}{\varepsilon^{2}},\frac{y}{\varepsilon}\right)\,{\rm d}y\,{\rm d}s,$
@@
(2.7) $w_{\varepsilon}:=f_{\varepsilon}-f_{0}-\varepsilon\omega\left(\frac{x}{\varepsilon}\right)\cdot S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right),$
where $f_{\varepsilon}$ and $f_{0}$ are solutions to (2.1) and (2.4) respectively. Note that $w_{\varepsilon}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$ and that $w_{\varepsilon}(0,\cdot)\equiv 0$. The study of a function defined similar to (2.7) is carried out in the quantitative analysis of periodic elliptic problems [16, Chapter 3, page 33]. We note, however, that the analysis carried out by Geng and Shen in [14, equation (1.14), page 2095] considers a function which also involves a term of order $\mathcal{O}(\varepsilon^{2})$. We will be employing such a construction in the next section. In the remainder of this manuscript, we will sometimes be using the following shorthand notations:
@@
(2.8) $\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$
@@
(2.9) $\operatorname{div}_{y}\phi(y)=\Theta(y)+\Theta(y)\nabla_{y}\omega(y)-\Theta_{h}\qquad\mbox{ in }Y.$
@@
(2.10) $\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$
@@
where we have used the fact that $f_{\varepsilon}$ and $f_{0}$ satisfy the evolution equations in (2.1) and (2.4), respectively. Owing to the anti-symmetric nature of the third-order tensor (i.e. $\phi_{kij}=-\phi_{ikj}$ for $1\leq i,j,k\leq d$), it follows that (see [16, Remark 3.1.3]) for a nice function $\psi$, we have
@@
(2.11) $\left|\int_{\Omega}\tau\left(\frac{x}{\varepsilon}\right)\kappa(x)\,{\rm d}x\right|\leq C\varepsilon\left\|\nabla\kappa\right\|_{\mathrm{L}^{\frac{2d}{d+2}}(\Omega)}.$
@@
(2.12) $\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi$
@@
where we have used the Lemma A.1 from the Appendix. Note that
@@
where we have used Lemma A.2 from the Appendix and have employed the Hölder inequality in the last step. Using a property of $S_{\varepsilon}$ (see Lemma A.3 from the Appendix), we arrive at the following bound:
@@
where we have used the fact [14, Remark 3.6, page 2106] that for any $g\in\mathrm{H}^{1}(\Omega)$,
@@
and the following estimate [14, Remark 3.10, page 2108]:
@@
(2.13) $\nabla\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}=S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)+S_{\varepsilon}\Big(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla^{2}f_{0}\Big).$
@@
where we have used Lemma A.1 from the Appendix. Using the property (2.13), and arguing as we did in the case of $I_{4}$, we can arrive at the following bound for $I_{5}$:
@@
3. The case of regular coefficients
@@
(3.1) $\left\|f_{\varepsilon}-f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{2}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}.$
@@
(3.2) $P:=\Theta(y)+\Theta(y)\nabla_{y}\omega(y)-\Theta_{h}\qquad\mbox{ in }Y.$
@@
(3.3) $H_{ij}:=\left\{\begin{aligned} P_{ij}&\qquad\mbox{ for }1\leq i,j\leq d,\\ -\zeta(y)\omega_{j}(y)&\qquad\mbox{ for }i=d+1\mbox{ and }1\leq j\leq d.\end{aligned}\right$.
@@
Hence there exists a third order tensor $\mathfrak{B}$ (see [16, Proposition 3.1.1, page 35]) such that
(3.4) $\frac{\partial\mathfrak{B}_{(d+1)ij}}{\partial\tau}+\sum_{k=1}^{d}\frac{\partial\mathfrak{B}_{kij}}{\partial y_{k}}=H_{ij}\qquad\mbox{ for }1\leq i\leq d+1,\quad 1\leq j\leq d,$
@@
(3.5) $\mathfrak{B}_{kij}=-\mathfrak{B}_{ikj}\qquad\mbox{ for }1\leq i,k\leq d+1,\qquad 1\leq j\leq d.$
@@
(3.6) $z_{\varepsilon}:=f_{\varepsilon}-f_{0}-\varepsilon\omega\left(\frac{x}{\varepsilon}\right)\cdot S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\nabla f_{0}\right)-\varepsilon^{2}\sum_{i,j=1}^{d}\frac{\mathfrak{B}_{(d+1)ij}\left(\frac{x}{\varepsilon}\right)}{\zeta\left(\frac{x}{\varepsilon}\right)}\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right),$
@@
(3.7) $\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}z_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla z_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$
@@
(3.8) $\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}z_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla z_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$
@@
(3.9) $\ell_{1}\leq C\varepsilon\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;\mathrm{W}^{1,d}(\Omega))}\left\|\psi\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}.$
@@
(3.10) $\ell_{2}\leq C\varepsilon^{\frac{1}{2}}\left\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{\text{in}}\right\|_{\mathrm{L}^{2}(\Omega)}\right\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$
@@
(3.11) $\ell_{3}\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$
@@
(3.12) $\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right)=S_{\varepsilon}\left(\frac{\partial}{\partial x_{i}}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\frac{\partial^{2}f_{0}}{\partial x_{i}\partial x_{j}}\right)$
@@
(3.13) $\nabla^{2}S_{\varepsilon}\left(\eta_{3}^{\varepsilon}\eta_{4}^{\varepsilon}\nabla f_{0}\right)=\nabla S_{\varepsilon}\left(\nabla\left(\eta^{\varepsilon}_{3}\eta^{\varepsilon}_{4}\right)\nabla f_{0}\right)+\nabla S_{\varepsilon}\left(\eta^{\varepsilon}_{3}\eta^{\varepsilon}_{4}\nabla^{2}f_{0}\right)$
@@
Using (3.12) and (3.13) and a few properties of the smoothing operator $S_{\varepsilon}$ (see Lemma A.1 and A.3 in the Appendix), we obtain the bound:
(3.14) $\ell_{4}$ $\leq C\|\nabla f_{0}\|_{\mathrm{L}^{2}(\Omega_{T,3\varepsilon})}\|\nabla\psi\|_{\mathrm{L}^{2}(\Omega_{T,3\varepsilon})}+C\varepsilon\|\nabla^{2}f_{0}\|_{\mathrm{L}^{2}(\Omega_{T})}\|\nabla\psi\|_{\mathrm{L}^{2}(\Omega_{T})}$
@@
Note that to arrive at the estimate (3.14), we have used the fact that $\frac{\mathfrak{B}_{(d+1)ij}}{\zeta}\in\mathrm{H}^{1}(\Omega)$ – see Lemma A.4 in the Appendix. We emphasise here that it is at this point we require the coefficient $\zeta$ to be more regular. The terms $\ell_{5}$ and $\ell_{6}$ are similar to $I_{5}$ in the proof of Lemma 2.4 (again, except for the choice of the cut-off functions) and hence can be handled in a similar fashion while taking into account the identity: